OpenStudy (anonymous):

a, 2a, 4a, 8a If a < 0, which of the four numbers above is the greatest? A. a B. 2a C. 4a D. 8a E. It cannot be determined from the information given. Thanks! My initial hunch is E. Is that right?

myininaya (myininaya):

E

myininaya (myininaya):

if a was an integer, we could say the answer is A

OpenStudy (anonymous):

The question did not state that a is an integer, so E would be the best answer?

myininaya (myininaya):

for instance let a=-1/2 2a=2*-1/2=-1 a>2a let's think some more let a=-5/2 then 2a=2*-5/2=-5 a>2a wait math can you think of a counterexample

myininaya (myininaya):

lets see if we can think of a counterexample

myininaya (myininaya):

math might be right

myininaya (myininaya):

I think A is correct now I can't think of a counterexample to what i said

myininaya (myininaya):

math say something so i can give you a medal

OpenStudy (anonymous):

Same here, I can't think of any either. Thanks Myiniaya!

OpenStudy (anonymous):

so the answer is a

myininaya (myininaya):

i think so i can't think of anything that proves you wrong

myininaya (myininaya):

i don't know what i was thinking i think i got this confused with a more confusing question that i seen before and i thought it was a trick

OpenStudy (anonymous):

@Myiniaya: Me too. Thanks so much to both of you! :D

myininaya (myininaya):

np